10 1 10 10 — X 30 30 1. Calculate the value of h for which lx = ly for the shaded area
Answer #1

...->n 30 30 : 2 = 0 y = A, g, + AzZ A+ A2 jó = Goxio xin + 20xhx ( +10) 60 X10 + 20xh 9 : 300 + 20k (6 + 10) _ 150 + (y) ScaBy = lo x +(5-)x10x60 + 5x20 + 2007 (335 * 1loyee? = { 150 40() Jžco + kepala 30th + 201 10 - 150 30th J - * te ay can, hogy